# The Internal Model Principle of Time-Varying Optimization

### v1

@@ II-A Problem statement
  3. (A3)
    The vector field $s(\theta)$ as well as $\theta(0)$ are known.
Under these assumptions, different methods have been developed in the literature, each with corresponding guarantees:
  1. (M1)
    Under (A1), all methods from static optimization (e.g., [23, 24, 19, 20, 21, 22, 13]) can be applied directly. All these methods can achieve convergence only to a neighborhood of a critical trajectory.
  2. (M2)
    The approach of [12] is applicable in this case, which is however limited to strongly convex quadratic optimization problems with linear temporal variablities.
  3. (M3)
    The class of prediction correction methods [10] and their variations [11] take advantage of this knowledge and ensures asymptotic tracking of $x^{\star}(t)$.
A natural question emerges directly from the above discussion: is (A3) necessary to ensure exact tracking and, if not, can it be relaxed (e.g., to (A1) or (A2))? With this motivation, in this paper we are driven by the following question.
Problem 0 (Minimal knowledge for exact tracking).
What is the least-restrictive set of assumptions (cf. (A1)–(A3)) under which there exists an optimization algorithm that achieves exact asymptotic tracking of the critical points of (1)?  $\square$
Our objective is to seek an optimization algorithm that assumes no access to $\theta(t)$; instead, as common in first-order optimization approaches [25], we will assume only the availability of functional evaluations of the gradient function $\nabla_{x}f(x,\theta(t))$ at points $x\in{}^{n},$ selected by the algorithm. Formally, the optimization algorithm will be described by an internal state $z(t),$ which takes values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$; the optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal) at which the gradient shall be evaluated, and processes functional evaluations of the gradient at these points, $y(t)=\nabla_{x}f(x(t),\theta(t))$ (called gradient feedback signal). Mathematically, the optimization algorithm is:
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (4a)
together with the gradient-feedback signal:
$y(t)=\nabla_{x}f(x(t),\theta(t)),$ (4b)
where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed. In the remainder, we refer to (4) as a dynamic gradient-feedback optimization algorithm. The architecture of the studied gradient-feedback algorithm is illustrated in Figure 1. Notice that the dynamics of the optimization algorithm (4), coupled with the time-variability generator (2), have the form of a nonlinear autonomous system:
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$
$y(t)$ $=\nabla_{x}f(G_{c}(z(t)),\theta(t)),$
$\dot{\theta}(t)$ $=s(\theta(t)).$ (5)
Our objective is to design $F_{c}(z,y),$ $G_{c}(z),$ and $n_{c},$ so that $y(t)\rightarrow 0$ as $t\rightarrow\infty,$ which ensures that $x(t)$ tracks, with zero asymptotic error, a critical trajectory $x^{\star}(t)$ of (1).
We will assume that $\theta(t)$ takes values in a neighborhood of the origin, and thus let $\Theta$ be some neighborhood of the origin of ${}^{p}.$ Note that there is no loss of generality in doing so, because if $\theta(t)$ takes values in the neighborhood of any other point, the former can be shifted to the origin via a change of variables without altering the critical points of (1). In what follows, we will denote by $x^{\star}_{\circ}\in{}^{n}$ a point such that
$0=\nabla_{x}f(x^{\star}_{\circ},0),$ (6)
and assume that $x^{\star}_{\circ}\in{}^{n}$ is locally unique. Moreover, we will let $z^{*}_{\circ}$ be such that $x_{\circ}^{\star}=G_{c}(z^{\star}_{\circ}),$ and assume that $z^{*}_{\circ}$ exists and is locally unique.
Definition 1.
We say that (4) asymptotically tracks a critical trajectory of (1) if, for each initial condition $(z(0),\theta(0))$ in some neighborhood of $(z_{\circ}^{*},0)\in\mathcal{Z}\times\Theta$, the solution of (II-A) satisfies $y(t)\to 0$ as $t\to\infty$. $\square$
Notice that, when an algorithm asymptotically tracks a critical trajectory, $x(t)\rightarrow x^{\star}(t)$ for some critical trajectory $x^{\star}(t);$ namely, critical point is readily given by the exploration signal. We make the objective of this work formal as follows.
Problem 1 (Dynamic gradient-feedback problem).
Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a gradient-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1). Derive a method to design such an algorithm when these conditions hold. $\square$
It is important to note that existence conditions for $F_{c}(z,y)$ and $G_{c}(z)$ shall not depend on the temporal variability of the optimization (namely, $s(\theta)$), but only on the properties of the optimization problem. Conversely, if such conditions were to depend on $s(\theta)$, a more general class of optimization algorithms than (4) could be constructed, implying the formulation (4) would not be sufficiently general.
II-B Standing assumptions
We list hereafter the basic assumptions on which our approach to gradient-feedback theory is based.
Assumption 3 (Stability of the exosystem).
The point $\theta=0$ is a stable equilibrium of (2), and there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of 00 with the property that each initial condition $\theta(0)\in\Theta_{\circ}$ is Poisson stable [^1] Recall that, for a nonlinear system of the form (2), an initial condition $\theta^{\circ}$ is said to be Poisson stable if the solution starting from $\theta^{\circ}$ at time $t$ of (2) (here denoted by $\theta_{t}\left(\theta^{\circ}\right)$)) is defined for all $t\in\mathbb{R}_{\geq 0}$ and, for every neighborhood $U$ of $\theta^{\circ}$ and each $T\in\mathbb{R}_{>0}$, there exists a time $t_{1}>T$ such that $\theta_{t_{1}}(\theta^{\circ})\in U$, and a time $t_{2}<-T$ such that $\theta_{t_{2}}(\theta^{\circ})\in U$.. $\square$
@@ II-B Standing assumptions
For convenience, we will require that $F_{c}(z^{\star}_{\circ},0)=0.$ This ensures that the optimization algorithm (4) has an equilibrium at $z=z^{\star}_{\circ},$ and the corresponding gradient feedback signal is identically zero: $y(t)=\nabla_{x}f(x_{\circ}^{\star},0)=0.$ Since $\theta(t)$ is unmeasurable, the temporal variability of the cost can only be evaluated through measurements of $y(t);$ hence, we make the following assumption.
Assumption 4 (Detectability of the exosystem).
Let
The pair $(Q,S)$ is detectable. $\square$
Detectability ensures that all the modes of the exosystem (2) can be reconstructed from the available measurements $y(t).$ Undetectability of $(Q,S)$ corresponds to a redundant description of the exogenous signal: indeed, if some modes of the exosystem do not influence the gradient, then $y(t)$ would be independent of those modes and they can thus be removed without altering the problem.
Example 1 (Detectability).
Suppose the objective function admits the decomposition $f(x,\theta(t))=\hat{f}(x)+x^{\mathsf{T}}\theta_{1}(t),$ where $\theta_{1}(t)=\alpha\cos(\omega t)$ for some $\alpha$ and $\omega$. The parameter signal can be generated by (2) with initial condition $\theta(0)=(\alpha,0)$ and vector field:
$s(\theta)=\begin{bmatrix}0&1\\ -\omega^{2}&0\end{bmatrix}\theta.$
Notice that, although the gradient signal $y(t)=\nabla_{x}\hat{f}(x)+\theta_{1}(t)$ is independent of $\theta_{2}(t),$ both modes of the exosystem affect the gradient (indeed, $\theta_{1}(t)=\alpha\cos(\omega t)$ cannot be generated by a linear system with a single mode). Notice also that the pair $(Q,S)$ is detectable. $\square$
III The parameter-feedback problem
To address Problem 1, we first study a simpler problem that allows us to derive the necessary framework to tackle our objectives in their generality. To this end, in place of the dynamic optimization algorithm (4), we begin by considering an algebraic optimization algorithm of the form:
@@ III The parameter-feedback problem
where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed; we will require that $H_{c}$ is of class $C^{0}$ and satisfies the fixed-point condition $x^{\star}_{\circ}=H_{c}(0).$ Because of the explicit dependence on $\theta(t),$ we will refer to (8) to as a parameter-feedback optimization algorithm. Our objective is to design the map $H_{c}$ so that the composition of (2), (4b), and (8):
$y(t)$ $=\nabla_{x}f(H_{c}(\theta(t)),\theta(t)),$
$\dot{\theta}(t)$ $=s(\theta(t)),$ (9)
tracks, with zero asymptotic error, the critical trajectories of (1). For the framework considered in this section, Problem 1 is reformulated as follows.
Problem 2 (Static parameter-feedback problem).
Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a parameter-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1). Derive a method to design such an algorithm when these conditions hold. $\square$
Solvability of the static parameter-feedback problem will depend on the existence of a function that zeros the gradient.
@@
Moreover, we say that $H_{c}$ zeros the gradient globally if (10) holds for all $\theta\in\Theta$, and locally if (10) holds for all $\theta\in\Theta_{\circ},$ where $\Theta_{\circ}\subseteq{}^{p}$ is some neighborhood of the origin [^2] The neighborhoods in Defn. 2 and Assumption 3 need not be the same.. $\square$
The following result characterizes all parameter-feedback optimization algorithms that achieve asymptotic tracking of a critical trajectory.
Theorem 1 (Parameter-feedback algorithm characterization).
Let Assumptions 1, 2, and 3 hold. The parameter-feedback algorithm (III) asymptotically tracks a critical trajectory of (1) if and only if the mapping $H_{c}$ zeros the gradient locally.
@@
(Only if) Suppose $y(t)\to 0$ as $t\to\infty$. We will show that $H_{c}$ zeros the gradient locally. By Poisson stability of the exosystem (cf. Assumption 3), there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that, for every $\theta_{\circ}\in\Theta_{\circ},$ every $\varepsilon>0,$ and every $T>0,$ the trajectory of (III) satisfies
$\|\theta(t)-\theta_{\circ}\|<\varepsilon,$
at some $t>T$. Then, $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (10) holds for all $\theta\in\Theta_{\circ}$.
(If) Suppose $H_{c}$ zeros the gradient on some neighborhood $\Theta_{0}\subset\Theta$ of the origin, i.e., (10) holds for all $\theta\in\Theta_{0}$. By Assumption 3, the point $\theta=0$ is a stable equilibrium of the exosystem, so there exists some other neighborhood $\Theta_{1}\subset\Theta$ of the origin and time $T>0$ such that $\theta(t)\in\Theta_{0}$ for all $t>T$ when $\theta(0)\in\Theta_{1}$. Since $H_{c}$ zeros the gradient on $\Theta_{0}$, this implies that $y(t)\to\infty$ as $t\to\infty$. ∎
Remark 2.
Poisson stability of the exosystem is needed for the gradient zeroing condition to be necessary for asymptotic tracking of a critical trajectory (but is not needed for sufficiency). This is because Poisson stability implies that all points sufficiently close to the origin are revisited infinitely often, so $H_{c}$ must zero the gradient on a neighborhood of the origin. More generally, $H_{c}$ must zero the gradient on all limit points of $\theta(t)$$\square$
The result provides a full characterization of all parameter-feedback algorithms that ensure exact asymptotic tracking. In words, $x=H_{c}(\theta)$ is a parameter-feedback algorithm if and only if it zeros the gradient everywhere in a neighborhood of the critical point. As a byproduct, the result also provides a necessary and sufficient condition for the solvability of the parameter-feedback problem: the problem is solvable if and only if the set of solutions to the system of equations $0=\nabla_{x}f(x,\theta)$ can be expressed, everywhere in a neighborhood of the origin of $\Theta$, as the graph of a function $x=H_{c}(\theta).$ Finally, the theorem provides an explicit form for the desired parameter-feedback algorithm (8): this is given by $x(t)=H_{c}(\theta(t));$ this algorithm ensures that, for all initial states $\theta(0)$ sufficiently close to the origin, $y(t)\rightarrow 0$ as $t\rightarrow\infty.$ In this sense, Theorem 1 provides a complete answer to Problem 2. We illustrate the applicability of the result and the necessity of the provided condition in the following example.
@@
Figure 2: Investigation of the condition (10). (Left) Loss function $f(x,\theta)$ studied in Example 2, plotted for $\theta=0.$ (Right) Gradient $\nabla f(x,\theta).$ The function $f(x,0)$ admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}}.$ At $x^{\star}_{\circ,1},$ condition (10) is not satisfied, since with an an upward shift of the graph of $\nabla_{x}f(x,0),$ $x^{\star}_{\circ,1}$ is no longer a critical point of $f(x,0).$ On the other hand, (10) holds for $x^{\star}_{\circ,2},$ since $x^{\star}_{\circ,2}$ varies continuously as $\theta$ is perturbed. See Example 2 for a discussion.
@@
On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily-small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig. 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (10). This graphical observation can be formalized with the aid of the implicit function theorem [32], as described next. Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow{}^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied. $\square$
Existence of a map $x=H_{c}(\theta)$ as in (10) can be ensured for general problems with the aid of the implicit function theorem [32], as illustrated next.

### v2

@@ 2.2 Problem statement
A second distinction can be made in terms of what knowledge concerning the objective function is available. In this case, several (non mutually-exclusive) options can be found in the literature (see, e.g., [48, 26, 10, 11, 34, 12]):
  1. (O1)
    The algorithm has access to oracle evaluations of (9).
  2. (O2)
    The analytical form of the gradient $\nabla_{x}f(x,\theta)$ is known.
  3. (O3)
    The sensitivity $\nabla_{x\theta}f(x,\theta)$ of the cost with respect to the parameter is known.
  4. (O4)
    The Hessian $\nabla_{xx}^{2}f(x,\theta)$ of the cost is known.
  5. (O5)
    The objective $f(x,\theta)$ is quadratic in both $x$ and $\theta$.
Table 1: Summary of required knowledge and known guarantees by various established approaches. The symbol ()($\checkmark$)( ✓ ) is used when (O1) follows from (E1) and (O2). See Remarks 4, 5, and 7 for discussions.
Approximate (E1) (E2) (O1) (O2) (O3) (O4) (O5)
solutions $\theta(t)$ $s(\theta)$ $(t,x)\mapsto\nabla_{x}f(x,\theta(t))$ $\nabla_{x}f(x,\theta)$ $\nabla_{x\theta}f(x,\theta)$ $\nabla_{xx}^{2}f(x,\theta)$ $f$ quadratic
Basic gradient flow [48, 26]
Control-based methods [12]
Prediction-correction [10, 11, 34] (✓)
This work (Section 3) (✓)
This work (Section 4)
These are not assumptions we impose a priori in this work, but rather widely used requirements in the literature [48, 26, 10, 11, 34, 12]. Various methods have been developed in the literature based on different combinations of these assumptions: for example, under (E1) and (O1), the basic gradient flow algorithm (7) converges (asymptotically) to a neighborhood of the critical points (see Remark 4). Under assumptions (E1), (E2), (O2), (O3), and (O4), the prediction-correction method (8) converges (asymptotically or in fixed-time) exactly to the optimizers (see Remark 5). Table 1 provides a comparison of the assumptions required by established methods and their positioning with respect to this work.
Remark 7 (Generality of the algorithm class).
The gradient flow algorithm (7) can be viewed as a special case of (10) with $F_{c}(z,y)=-\eta y$ and $G_{c}(z,y)=z$. Similarly, the prediction-correction method (8) can be viewed as a special case of (10) with $z=(z_{1},z_{2})$ and
@@
$G_{c}(z)$ $=z_{1},$ (13)
with $z_{2}(0)=\theta(0).$ It follows that (10) is general enough to encompass these two established methods as special cases, and thus all conclusions concerning the class of algorithms (10) made here will also apply to these methods.  $\square$
Our second objective is to design optimization algorithms from the class (10) that track, with zero asymptotic error, the critical trajectories of (1). We formalize this notion next.
Problem 1 (Design of dynamic gradient-feedback optimization algorithms).
Design $F_{c}(z,y),$ $G_{c}(z),$ and $n_{c},$ requiring the minimal necessary knowledge for exact asymptotic tracking (as in Problem 2.2), so that (10) exactly asymptotically tracks a critical trajectory of (1).  $\square$
3 The parameter-feedback problem
In many cases of interest [26], having access to $y(t)$ is a byproduct of having access to both the function $\nabla_{x}f(x,\theta)$ and knowledge or measurements of $\theta(t)$; see Remark 6. In these cases, assumptions (E1) and (O2) are automatically satisfied. In this section, we analyze this scenario. We anticipate that the results derived in this section will also serve as an intermediate step to tackle the more challenging problem where assumption (E1) is relaxed, which is the focus of Section 4.
When the algorithm has access to both $\theta(t)$ at each time $t\in\mathbb{R}_{\geq 0},$ and the function $\nabla_{x}f(x,\theta)$, the measurements $y(t)$ do not provide additional information. Hence, under the assumptions of this section, the optimization algorithm (10) can be replaced by the algebraic relationship [^5] While one could consider a dynamic optimization algorithm of the form $\dot{z}(t)=F_{c}(z(t),\theta(t))$ and $x(t)=G_{c}(z(t)),$ we will prove in Theorem 1 that a dynamic structure is unnecessary.
@@ 3 The parameter-feedback problem
where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed. Because of the explicit dependence on $\theta(t),$ we will refer to (14) to as a parameter-feedback optimization algorithm. In analogy with (11), for simplicity of the presentation, in this section, we will impose that $H_{c}(\theta)$ satisfies:
$x^{\star}_{\circ}=H_{c}(0),$
and require that $H_{c}$ is of class $C^{0}.$
The composition of (2), (10a), and (14) is given by:
$y(t)$ $=\nabla_{x}f(H_{c}(\theta(t)),\theta(t)),$ (15a)
$\dot{\theta}(t)$ $=s(\theta(t)).$ (15b)
In analogy with Definition 1, we will say that the parameter-feedback algorithm (14) exactly asymptotically tracks a critical trajectory of (1) if there exists a neighborhood $\Theta_{s}\subset\Theta$ of the origin such that, for each initial condition $\theta(0)\in\Theta_{s}$, the solution of (15) satisfies $y(t)\to 0$ as $t\to\infty$.
For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated as follows.
Problem 2 (Minimal knowledge for parameter feedback).
Consider the class of optimization algorithms (14). Determine the minimal necessary knowledge concerning (1) and (2), needed to design an algorithm from this class that exactly asymptotically tracks a critical trajectory of (1). $\square$
Problem 3 (Design of parameter-feedback optimization algorithms).
Design $H_{c}(\theta),$ requiring minimal knowledge for parameter-feedback (as in Problem 2), so that (14) exactly asymptotically tracks a critical trajectory of (1).  $\square$
3.1 Fundamental results
Solvability of the parameter-feedback problem will depend on the existence of a function that zeros the gradient on the set of limit points of the exosystem; we now define these notions.
@@ 3.1 Fundamental results
Moreover, we say that $H_{c}$ zeros the gradient on a set $\Theta_{\circ}\subseteq\Theta$ if (16) holds for all $\theta\in\Theta_{\circ}$. $\square$
Definition 3 (Limit point and limit set).
A point $\theta_{\omega}\in\Theta$ is a limit point with respect to the initialization $\theta_{\circ}\in\Theta$ if there exists a sequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ with $t_{i}\to\infty$ as ${i\to\infty}$ such that the exosystem (2) with $\theta(0)=\theta_{\circ}$ satisfies $\theta({t_{i}})\to\theta_{\omega}$ as ${i\to\infty}$. Let $\Omega(\theta_{\circ})$ denote the set of all limit points (i.e., for all sequences $t_{i}$) of (2) with respect to the initialization $\theta_{\circ}\in\Theta.$ Given $\Theta_{\circ}\subseteq\Theta,$ the set $\Omega(\Theta_{\circ}):=\cup_{\theta_{\circ}\in\Theta_{\circ}}\Omega(\theta_{\circ})$ is called the limit set with respect to initializations in $\Theta_{\circ}$ [49$\square$
Intuitively, $\Omega(\Theta_{\circ})$ denotes the set of all limit points (equilibria, limit cycles, etc.) that can be reached by the exosystem when initialized at points in $\Theta_{\circ}$. By the boundedness trajectories assumption (Assumption 2), $\Omega(\Theta_{\circ})$ is a bounded set.
The following result characterizes all parameter-feedback optimization algorithms that achieve asymptotic tracking.
Theorem 1 (Existence and characterization of parameter-feedback algorithms).
Let Assumptions 12 hold. The parameter-feedback algorithm (14) asymptotically tracks a critical trajectory of (1) if and only if there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that the mapping $H_{c}$ zeros the gradient on the limit set $\Omega(\Theta_{\circ})$.
@@
(Only if) Suppose $y(t)\to 0$ as $t\to\infty$ for initializations $\theta(0)\in\Theta_{\circ}$; we will show that $H_{c}$ zeros the gradient on $\Omega(\Theta_{\circ})$. By Assumption 2, the trajectories of the exosystem are bounded, and thus, by the Bolzano–Weierstrass theorem, there exists an increasing subsequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ such that $\theta(t_{i})$ converges to some limit point $\theta_{\omega}\in\Omega\left(\theta(0)\right)$. We then have:
$\lim_{i\to\infty}y({t_{i}})=\lim_{i\to\infty}\nabla_{x}f(H_{c}(\theta({t_{i}})),\theta({t_{i}}))=\nabla_{x}f(H_{c}(\theta_{\omega}),\theta_{\omega}),$ (17)
where the second identity follows by the continuity of the gradient (see Assumption 1) and that of $H_{c}$. Because, $y(t)\to 0$, the left-hand side of (17) is equal to zero, and thus (17) implies that $\nabla_{x}f(H_{c}(\theta_{\omega}),\theta_{\omega})=0.$ Since this holds for any limit point $\theta_{\omega}\in\Omega(\Theta_{\circ})$, the statement follows.
(If) Suppose $\theta(0)\in\Theta_{\circ}$ and that $H_{c}$ zeros the gradient on $\Omega(\Theta_{\circ})$. Then, the right-hand side of (17) is equal to zero, which implies the existence of a sequence $t_{i}$ such that $y({t_{i}})\to 0$ as ${i\to\infty}$. Since this holds for any limit point $\theta_{\omega}\in\Omega(\theta(0))$, any convergent subsequence of $y(t)$ converges to zero. Moreover, $y(t)$ is bounded due to the Lipschitz continuity of the gradient, so $y(t)$ also converges to zero as $t\to\infty$. By iterating the reasoning for all $\theta(0)\in\Theta_{\circ},$ it follows that $y(t)\to 0$ for all $\theta(0)\in\Theta_{\circ}$, and the claim follows. ∎
The theorem shows that the existence of a parameter-feedback algorithm is dependent upon the existence of a mapping that zeros the gradient. An application of this theorem is illustrated in Example 1, which shows that a parameter-feedback algorithm may fail to exist in some circumstances. The theorem also provides a full characterization of all parameter-feedback algorithms that achieve asymptotic tracking. Loosely speaking, $x=H_{c}(\theta)$ is a parameter-feedback algorithm if and only if it zeros the gradient on the limit set of the exosystem. As a byproduct, the result also provides a necessary and sufficient condition for the solvability of the parameter-feedback problem: the problem is solvable if and only if the set of solutions to the system of equations $0=\nabla_{x}f(x,\theta)$ can be expressed, at all limit points of the exosystem, as the graph of a function $x=H_{c}(\theta).$ We present a set of sufficient conditions for this to hold in Section 3.2.
By Theorem 1, the problem of designing a parameter-feedback algorithm for exact asymptotic tracking can be reduced to that of finding a mapping $x=H_{c}(\theta)$ that zeros the gradient. Therefore, we have the following.
Answer to Problem 2. When a parameter-feedback algorithm that achieves exact tracking exists, the minimal knowledge needed to design such an algorithm is (O2). Notice that assumption (E1) is implicitly required to execute the algorithm (14).
We illustrate the applicability of the result and the necessity of the provided condition in the following example.
@@
Figure 2: Investigation of the condition (16). (Left) Loss function $f(x,\theta)$ studied in Example 1, plotted for $\theta=0.$ (Right) Gradient $\nabla f(x,\theta).$ The function $f(x,0)$ admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}}.$ At $x^{\star}_{\circ,1},$ condition (16) is not satisfied, since with an upward shift of the graph of $\nabla_{x}f(x,0),$ $x^{\star}_{\circ,1}$ is no longer a critical point of $f(x,0).$ On the other hand, (16) holds for $x^{\star}_{\circ,2},$ since $x^{\star}_{\circ,2}$ varies continuously as $\theta$ is perturbed. See Example 1 for a discussion.
@@
On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig. 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (16). This graphical observation can be formalized with the aid of the implicit function theorem [47], as described next. Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow{}^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied. $\square$
We conclude by observing that the proof of Theorem 1 is constructive in that it provides an explicit procedure to construct parameter-feedback algorithms.
Answer to Problem 3. When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge (O2) is to first determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$ and then find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set, in which case the parameter-feedback algorithm is given by (14).
We conclude by illustrating the parameter-feedback design procedure on a quadratic problem.
Example 2.
Consider an instance of (1) with quadratic cost and time-variability that depends linearly on $\theta(t)$ (which, e.g., has been investigated in [12]):
with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
$y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$
For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique. Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin. Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$. Note that, by substituting into (15), we have ${y(t)=RH_{c}\theta(t)+Q\theta(t)=0}$ for all times $t\in\mathbb{R}_{\geq 0}.$ Namely, the gradient is identically zero. This property holds because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin). In fact, this feature holds for the any optimization problem for which the gradient of the cost admits a mapping that zeros the gradient globally. A set of sufficient conditions to check for this property are given in Theorem 3. $\square$
Remark 8 (Prediction-correction algorithms asymptotically compute maps that zero the gradient).
As an illustration, consider the quadratic problem studied in Example 2. Let $x_{\text{PC}}(t)$ denote the iterates of the prediction-correction algorithm (8) applied to the quadratic objective (19), and let $x_{\text{PF}}(t)=H_{c}\theta(t)$ with $H_{c}=-R^{\dagger}Q$ be the parameter-feedback algorithm. Then, the error $e(t)=x_{\text{PC}}(t)-x_{\text{PF}}(t)$ satisfies the dynamics $\dot{e}(t)=-e(t)$, indicating that the iterates of prediction-correction converge exponentially to those of the parameter-feedback algorithm, which zeros the gradient. $\square$
3.2 Sufficient conditions for the existence of a parameter-feedback algorithm
Existence of a map $x=H_{c}(\theta)$ as in (16) can be ensured for general problems with the aid of the implicit function theorem [47], as illustrated by the following result.